OpenStudy (anonymous):

What is the surface area of a sphere with a radius of 11 centimeters

OpenStudy (amistre64):

is the formula; 4/3 pi r^3 ?

OpenStudy (amistre64):

no .. its akin to circumferent

OpenStudy (anonymous):

do 3.14x11

OpenStudy (amistre64):

4 pi r maybe

OpenStudy (amistre64):

4pi r^2 .... that would be it

OpenStudy (anonymous):

3.14x11

OpenStudy (amistre64):

http://www.math.com/tables/geometry/surfareas.htm

OpenStudy (zarkon):

\[\frac{d}{dr}\left(\frac{4}{3}\pi r^3\right)=4\pi r^2\] :)

OpenStudy (amistre64):

\[2*\int_{0}^{r} 2\pi\ r\ dr\]

OpenStudy (amistre64):

lol ... that too